# Generalized Rate Operator Quantum Jumps via Realization-Dependent Transformations

### v1

@@ Abstract
The dynamics of open quantum systems is often solved by stochastic unravellings where the average over the state vector realizations reproduces the density matrix evolution. We focus on quantum jump descriptions based on the rate operator formalism. In addition to displaying and exploiting different equivalent ways of writing the master equation, we introduce state-dependent rate operator transformations within the framework of stochastic pure state realizations, allowing us to extend and generalize the previously developed formalism. As a consequence, this improves the controllability of the stochastic realizations and subsequently greatly benefits when searching for optimal simulation schemes to solve open system dynamics. At a fundamental level, intriguingly, our results show that it is possible to have positive unravellings – without reverse quantum jumps and avoiding the use of auxiliary degrees freedom – in a number of example cases even when the corresponding dynamical map breaks the property of P-divisibility, thus being in the strongly non-Markovian regime.
@@ I Introduction
The MCWF method has been generalized to non-positive rates by the non-Markovian quantum jumps (NMQJ) method [29, 30]. However, the different stochastic realizations are no longer independent, thus making the simulations more expensive. Nevertheless, it was shown that it is possible to generalize the MCWF method and maintaining independent realization also under the weaker assumption of positive(P)-divisibility of the dynamical maps [31]. This method relies on the definition of the rate operator (RO) [32, 33, 34, 35] and was therefore named rate operator quantum jumps (ROQJ). Violations of P-divisibility have also been linked to a different definition of non-Markovianity [26, 36]. Interestingly, these unravelings relies on jumps to mutually orthogonal states and have been linked to the study of pointer bases [37, 38].
Recently, in [39], the RO formalism has been expanded by employing the non-unique ways of writing the master equation, thus leading to the formulation of a family of distinct RO unravelings for the same master equation. In this paper, we generalize those results by allowing the RO to depend on the current state of the particular realization. This generalized RO not only enhances the efficiency of controlling the stochastic realizations, but also enables the characterization of certain dynamics violating P-divisibility, thus non-Markovian according to all definitions, while preserving the independence between the different realizations. Noticeably, we are able to do so without requiring any additional ancillary degree of freedom.
@@ II.1 Open quantum systems
for all orthonormal basis $\{\varphi_{i}\}_{i}$, for all $i\neq j$.
@@
P- but not CP-divisible $\times$ $\circ *{}^{*}$
@@
Table 1: Comparison of the $\psi$-ROQJ unravelings with the previously introduced methods; ✓: yes, $\times$: no, $\circ$: sometimes, $*{}^{*}$: yes, unless the some rates are negative since the beginning of the dynamics. Each column refers to a distinct unraveling method, the first three rows denote the applicability of the methods to different classes of dynamics, while the fourth row denotes the possibility to realize each trajectory independently from the others. We note in particular that $\psi$-ROQJ is the only method that can tackle also some non-P-divisible-dynamics, while keeping the different realizations independent, as shown in Sec. V.2. As explained in the text, the $\psi$-ROQJ includes the $W$- and the $R$-ROQJ as special cases, while there is no inclusion among the latter two. The table refers to ROQJ methods without reversed jumps; note that all the ROQJ methods can be supplemented with reverse jumps [29], leading to correlated realizations and extending their range of applicability to non-P-divisible dynamics.
@@ II.2 Quantum jump unravelings
In the literature, there have been introduced many different unravelings consisting of piecewise deterministic processes on the set of pure states on the system’s Hilbert space $\mathscr{H}$. The exact dynamics of Eq. (1) is reproduced by averaging over all stochastic realizations, with the form of the deterministic and jump process that can vary significantly for the different unraveling methods. Differently from other methods, our approach does not use additional degree of freedom [46, 47, 48, 49, 50, 51], which would require additional computational effort, nor temporarily negative probabilities for the occupation of certain states [52].
@@
$p^{\alpha}_{\psi(t),t}=\gamma_{\alpha}(t)\lVert L_{\alpha}(t)\ket{\psi(t)}\rVert^{2}dt.$ (10)
@@
$W_{\psi}\coloneqq\big{(}\mathbbm{1}-P_{\psi}\big{)}\mathcal{J}_{t}[P_{\psi}]\big{(}\mathbbm{1}-P_{\psi}\big{)},$ (11)
where $P_{\psi}=\ket{\psi}\bra{\psi}$ |, with probabilities given by the corresponding eigenvalues multiplied by the infinitesimal time increment $dt$. Following the nomenclature of [39], we call this unraveling method W-ROQJ, emphasising that the jumps and their probabilities are fixed by the eigenstates and eigenvalues of $W$. We further note that the pre-jump state $\psi$ is an eigenstate of $W_{\psi}$, so that the state after the jump is always orthogonal to $\psi$. The deterministic evolution is generated by the non-Hermitian non-linear effective Hamiltonian $K^{W}_{\psi,t}=K(t)+\Delta_{\psi,t}$, with
$\Delta_{\psi,t}=\frac{i}{2}\sum_{\alpha}\gamma_{\alpha}(t)\left(2L_{\alpha}(t)\ell^{*}_{\psi,\alpha}(t)-\lvert\ell^{*}_{\psi,\alpha}(t)\rvert^{2}\right),$ (12)
where $\ell_{\psi,\alpha}(t)=\braket{\psi}{L_{\alpha}(t)}{\psi}$. The operator $W_{\psi}$ does not depend on the particular form (6) of the master equation [33, 34] and is positive for all states $\psi$ if and only if the dynamics is P-divisible [10]. Therefore, such method can be used to unravel any P-divisible dynamics, significantly beyond the range of applicability of the MCWF.
The ROQJ method can be extended [39] relying on the non-uniqueness of the decomposition of the master equation (6)-(7) and considering jumps |$\ket{\psi(t)}\mapsto\ket{\varphi_{\psi(t),t}^{j}}$ to the eigenstates of the RO
$R_{\psi(t),t}\coloneqq\mathcal{J}_{t}^{\prime}[P_{\psi(t)}]=\mathcal{J}_{t}[P_{\psi(t)}]+\frac{1}{2}(C(t)P_{\psi(t)}+P_{\psi(t)}C^{\dagger}(t))$ (13)
@@
Going beyond the MCWF, it is possible to deal with temporarily negative rates by using the NMQJ technique as in [29, 30], by considering reverse jumps |$\ket{\psi_{i}(t)}=L_{\alpha}(t)\ket{\psi_{j}(t)}/\lVert L_{\alpha}(t)\ket{\psi_{j}(t)}\rVert\mapsto\ket{\psi_{j}(t)}$, with probability depending on the rate $N_{j}(t)/N_{i}(t)$ between the occupations of |$\ket{\psi_{j}(t)}$ and |$\ket{\psi_{i}(t)}$. The same method can be employed similarly also for the RO whenever some eigenvalues are negative. As a drawback, using this method the different stochastic realizations become dependent one on the other, thus making the simulation more expensive. If, instead, the rates of the RO, and therefore the jump probabilities, are positive for all realizations, we say that such unraveling is a positive unraveling.
@@ III Generalized Rate Operator
Suppose that, at some time $t$, the state can be written in terms of the single realizations as $\rho=\sum_{i}p_{i}(t)\ket{\psi_{i}(t)}\bra{\psi_{i}(t)}$ |. From the point of view of a single realization |$\ket{\psi}$, it is possible to choose the transformation $C$ of Eq. (6)-(7) depending on the current state |$\ket{\psi}$ and leaving the average evolution unaffected. The new generalized RO is therefore of the form
$R_{\psi,t}\coloneqq\mathcal{J}_{t}[P_{\psi}]+\frac{1}{2}\big{(}C_{\psi,t}P_{\psi}+P_{\psi}C_{\psi,t}^{\dagger}\big{)},$ (15)
with $C_{\psi,t}$ that can depend non-trivially on the current state of the realization $\psi$. To emphasize the dependence on the state $\psi$, we refer to the generalized RO as $\psi$-ROQJ. Indeed, whenever $C_{\psi,t}$ does not depend on $\psi$, the $\psi$-ROQJ reduces to the $R$-ROQJ of Eq. (13). The deterministic evolution is non-linear because of the state-dependence of the effective non-Hermitian Hamiltonian
$K_{\psi,t}\coloneqq H(t)-\frac{i}{2}\Gamma(t)-\frac{i}{2}C_{\psi}(t).$ (16)
Since $C_{\psi}(t)$ appears only to apply to |$\ket{\psi}$, it is possible to simplify the transformation by defining the unnormalized vector
|$\ket{\Phi_{\psi,t}}\coloneqq C_{\psi,t}\ket{\psi}.$ (17)
@@
[…]
The unraveling is obtained by considering jumps |$\ket{\psi(t)}\mapsto\ket{\varphi_{\psi(t),t}^{j}}$ to the eigenstates of the RO, with probability $p^{j}_{\psi(t),t}=\lambda_{\psi(t),t}^{j}\,dt$, where $\lambda_{\psi(t),t}^{j}$ is the corresponding eigenvalue. The deterministic evolution is, up to normalization, given by
|$\begin{split}\ket{\psi(t)}\mapsto&\ket{\tilde{\psi}^{\text{det}}(t+dt)}=(\mathbbm{1}-iK_{\psi(t),t}\,dt)\ket{\psi(t)}\\ =&(\mathbbm{1}-iK(t)\,dt)\ket{\psi(t)}-\frac{dt}{2}\,\ket{\Phi_{\psi(t),t}}.\end{split}$ (19)
In Appendix A, we show that averaging over all the realizations indeed reproduces the master equation (1). The key argument is that each state |$\ket{\psi_{i}(t)}$ evolves, on average, as it would according to (1), for any possible choice of $R_{\psi_{i}(t),t}$.
@@ IV Characterization and limitations
Figure 1: Examples of the $\mathscr{H}_{\mathcal{J}_{t}}$ (red) and of $\Lambda_{t}(\mathcal{S}(\mathscr{H}))$ (green) for three non-P-divisible qubit dynamics at some fixed time $t$, showing the slice $y=0$ of the Bloch sphere. Left: no positive unraveling, since $\mathscr{H}_{\mathcal{J}_{t}}$ doesn’t contain any orthonormal basis. Middle: no positive unraveling, since $\Lambda_{t}\left(\mathcal{S}(\mathscr{H})\right)\not\subseteq\mathcal{P}(\mathscr{H}_{\mathcal{J}_{t}})$. Right: both necessary conditions hold, therefore a positive unraveling could exist.
@@ IV.1 Comparison with the $W$-ROQJ unravelings
By using the formalism of the $W$-ROQJ of Eq. (11), it is always possible to unravel any P-divisible dynamics by considering orthogonal jumps to the eigenstates of such operator. It is possible to obtain the same unravelings also with the $\psi$-ROQJ formalism by imposing that $\psi$ is an eigenstate of the RO: $R_{\psi}\ket{\psi}=\lambda^{\text{det}}\ket{\psi}$ for all $\psi$, where $\lambda^{\text{det}}$ is the corresponding eigenvalue. This way, the other post-jump states |$\ket{\varphi^{j}_{\psi}}$ are orthogonal to |$\ket{\psi}$ and it can be shown that the corresponding eigenvalues are (see Appendix B)
$\lambda^{j}_{\psi}=\sum_{\alpha}\gamma_{\alpha}\lvert\braket{\varphi^{j}_{\psi}}{L_{\alpha}}{\psi}\rvert^{2},$ (20)
@@ IV.2 Necessary conditions for a positive unraveling
We now investigate when it possible to have a positive unraveling for dynamics breaking P-divisibility. This possibility drastically simplifies the simulations since, for positive unravelings, the different realizations do not depend on each other. For any non-P-divisible dynamics, there always exists some state $\psi$ and time $t$ such that $R_{\psi,t}\not\geq 0$. In fact, from the condition for P-divisibility (2), there always exists a state $\psi_{\perp}$, orthogonal to $\psi$, such that
$\braket{\psi_{\perp}}{R_{\psi,t}}{\psi_{\perp}}=\sum_{\alpha}\gamma_{\alpha}(t)\lvert\braket{\psi_{\perp}}{L_{\alpha}(t)}{\psi}\rvert^{2}<0.$ (21)
Notice that this condition does not depend on the particular transformation $\Phi_{\psi,t}$ present in (18), but only on $\mathcal{J}_{t}$. This fact, however, doesn’t necessarily limit the existence of a positive unraveling: it could still exist if one is able to describe the state $\rho$ only using states for which $R_{\psi,t}\geq 0$.
@@
where $\mathscr{H}$ is the system Hilbert space. From Eq. (2), it is evident that a dynamics is P-divisible iff $\mathscr{H}_{\mathcal{J}_{t}}=\mathscr{H}$ $\forall t$. At variance with the the positivity domain introduced e.g. in [53, 54], which refers to a subset of the set $\mathcal{L}(\mathscr{H})$ of linear operators on $\mathscr{H}$, the positivity domain defined in Eq. (22) is a subset of $\mathscr{H}$.
A positive unraveling can exist only if it is possible to write any state as a convex combination $\rho(t)=\sum_{i}p_{i}(t)\ket{\psi_{i}(t)}\bra{\psi_{i}(t)}$ | using only states $\psi_{i}(t)\in\mathscr{H}_{\mathcal{J}_{t}}$ $\forall t$. Naturally, if $\mathscr{H}_{\mathcal{J}_{t}}=\emptyset$ for some time, it is not possible to have a positive unraveling. If we further assume invertibility of the dynamical map, then it is also not possible if $\mathscr{H}_{\mathcal{J}_{t}}$ is zero-measured: if, for any initial state, all stochastic realizations were inside a zero-measure set, then the whole set of quantum states $\mathcal{S}(\mathscr{H})$ would be mapped to a zero-measure set, thus breaking invertibility.
@@
These condition are depicted pictorially in Fig. 1. Interestingly, they do not depend on the unraveling but only on the dynamics $\Lambda_{t}$. Therefore it is possible to rule out the possibility to have a positive unraveling for some dynamics by simply looking at $\Lambda_{t}$. On the other hand, as we will see in Sec. V.3 such conditions can be satisfied even by non P-divisible dynamics, which in fact do admit a positive unraveling

### v2

@@ Abstract
The dynamics of open quantum systems is often solved by stochastic unravelings where the average over the state vector realizations reproduces the density matrix evolution. We focus on quantum jump descriptions based on the rate operator formalism. In addition to displaying and exploiting different equivalent ways of writing the master equation, we introduce state-dependent rate operator transformations within the framework of stochastic pure state realizations, allowing us to extend and generalize the previously developed formalism. As a consequence, this improves the controllability of the stochastic realizations and subsequently greatly benefits when searching for optimal simulation schemes to solve open system dynamics. At a fundamental level, intriguingly, our results show that it is possible to have positive unravelings – without reverse quantum jumps and avoiding the use of auxiliary degrees freedom – in a number of example cases even when the corresponding dynamical map breaks the property of P-divisibility, thus being in the strongly non-Markovian regime.
@@ I Introduction
The MCWF method has been generalized to non-positive rates by the non-Markovian quantum jumps (NMQJ) method [29, 30]. However, the different stochastic realizations are no longer independent, thus making the simulations more expensive. Nevertheless, it was shown that it is possible to generalize the MCWF method and maintaining independent realization also under the weaker assumption of positive(P)-divisibility of the dynamical maps [31]. This method relies on the definition of the rate operator (RO) [32, 33, 34, 35] and was therefore named rate operator quantum jumps (ROQJ). Violations of P-divisibility have also been linked to a different definition of non-Markovianity [26, 36]. Interestingly, these unravelings rely on jumps to mutually orthogonal states and have been linked to the study of pointer bases [37, 38].
Recently, in [39], the RO formalism has been expanded by employing the non-unique ways of writing the master equation by applying arbitrary time-dependent transformations, thus leading to the formulation of a family of distinct RO unravelings for the same master equation. In this paper, we generalize those results by allowing the arbitrary transformation to depend on the current state of the particular realization. This generalized RO not only enhances the efficiency of controlling the stochastic realizations, but also enables the characterization of certain dynamics violating P-divisibility, thus non-Markovian according to all definitions, while preserving the independence between the different realizations. Noticeably, we are able to do so without requiring any additional ancillary degree of freedom.
@@ II.1 Open quantum systems
for all orthonormal bases $\{\varphi_{i}\}_{i}$, for all $i\neq j$.
@@
P- but not CP-divisible $\times$ $\circ$
@@
Table 1: Comparison of the $\Psi$-ROQJ unravelings with the previously introduced methods; ✓: yes, $\times$: no, $\circ$: sometimes, : yes, unless some rates are negative since the beginning of the dynamics. Each column refers to a distinct unraveling method, the first three rows denote the applicability of the methods to different classes of dynamics, while the fourth row denotes the possibility to realize each trajectory independently from the others. We note in particular that $\Psi$-ROQJ is the only method that can tackle also some non-P-divisible-dynamics, while keeping the different realizations independent, as shown in Sec. V.2. As explained in the text, the $\Psi$-ROQJ includes the $W$- and the $R$-ROQJ as special cases, while there is no inclusion among the latter two. The table refers to ROQJ methods without reversed jumps; note that all the ROQJ methods can be supplemented with reverse jumps [29], leading to correlated realizations and extending their range of applicability to non-P-divisible dynamics.
@@ II.2 Quantum jump unravelings
In the literature, there have been introduced many different unravelings consisting of piecewise deterministic processes on the set of pure states on the system’s Hilbert space $\mathscr{H}$. The exact dynamics of Eq. (1) is reproduced by averaging over all stochastic realizations, with the form of the deterministic and jump process that can vary significantly for the different unraveling methods. Differently from other methods, our approach does not use additional degrees of freedom [46, 47, 48, 49, 50, 51], which would require additional computational effort, nor temporarily negative probabilities for the occupation of certain states [52].
@@
${p^{\alpha}_{\psi(t)}}=\gamma_{\alpha}(t)\lVert L_{\alpha}(t)\ket{\psi(t)}\rVert^{2}dt.$ (10)
@@
${W_{\psi(t)}\coloneqq\big{(}\mathbbm{1}-P_{\psi(t)}\big{)}\mathcal{J}_{t}[P_{\psi(t)}]\big{(}\mathbbm{1}-P_{\psi(t)}\big{)},}$ (11)
where $P_{\psi}=\ket{\psi}\bra{\psi}$ |, with probabilities given by the corresponding eigenvalues multiplied by the infinitesimal time increment $dt$. Following the nomenclature of [39], we call this unraveling method W-ROQJ, emphasising that the jumps and their probabilities are fixed by the eigenstates and eigenvalues of $W$. We further note that the pre-jump state $\psi(t)$ is an eigenstate of $W_{\psi(t)}$, so that the state after the jump is always orthogonal to $\psi(t)$. The deterministic evolution is generated by the non-Hermitian non-linear effective Hamiltonian $K^{W}_{\psi(t)}=K(t)+\Delta_{\psi(t)}$, with
${\Delta_{\psi(t)}=\frac{i}{2}\sum_{\alpha}\gamma_{\alpha}(t)\left(2L_{\alpha}(t)\ell^{*}_{\psi(t),\alpha}(t)-\lvert\ell^{*}_{\psi(t),\alpha}(t)\rvert^{2}\right),}$ (12)
where $\ell_{\psi(t),\alpha}(t)=\braket{\psi(t)}{L_{\alpha}(t)}{\psi(t)}$. The operator $W_{\psi(t)}$ does not depend on the particular form (6) of the master equation [33, 34] and is positive for all states $\psi$ if and only if the dynamics is P-divisible [10]. Therefore, such method can be used to unravel any P-divisible dynamics, significantly beyond the range of applicability of the MCWF.
The ROQJ method can be extended [39] relying on the non-uniqueness of the decomposition of the master equation (6)-(7) and considering jumps |$\ket{\psi(t)}\mapsto\ket{{\varphi_{\psi(t)}^{j}}}$ to the eigenstates of the RO
${R_{\psi(t)}\coloneqq\mathcal{J}_{t}^{\prime}[P_{\psi(t)}]=\mathcal{J}_{t}[P_{\psi(t)}]+\frac{1}{2}(C(t)P_{\psi(t)}+P_{\psi(t)}C^{\dagger}(t))}$ (13)
@@
Going beyond the MCWF, it is possible to deal with temporarily negative rates by using the NMQJ technique as in [29, 30], by considering reverse jumps |$\ket{\psi_{i}(t)}=L_{\alpha}(t)\ket{\psi_{j}(t)}/\lVert L_{\alpha}(t)\ket{\psi_{j}(t)}\rVert\mapsto\ket{\psi_{j}(t)}$, with probability depending on the ratio $N_{j}(t)/N_{i}(t)$ between the occupations of |$\ket{\psi_{j}(t)}$ and |$\ket{\psi_{i}(t)}$. However, one needs to know the average state $\rho(t)$ to compute the ratios, and therefore the realizations become dependent one on the other, making the simulation more expensive. The same method can be employed similarly also for the RO whenever some eigenvalues are negative. If, instead, the eigenvalues of the RO, and therefore the jump probabilities, are positive for all realizations, we say that such unraveling is a positive unraveling. In this case, the stochastic realizations are independent, since the jump probabilities can be calculated from the state of the given individual realization, and the simulation is more efficient.
@@ III Generalized Rate Operator
Suppose that, at some time $t$, the state can be written in terms of the single realizations as $\rho=\sum_{i}p_{i}(t){P_{\psi_{i}(t)}}$. From the point of view of a single realization |$\ket{\psi(t)}$, it is possible to choose the transformation $C$ of Eq. (6)-(7) to depend not only on time but also on the current state |$\ket{\psi(t)}$ and leaving the average evolution unaffected. The new generalized RO is therefore of the form
${\Psi\text{-}R_{\psi(t)}\coloneqq\mathcal{J}_{t}[P_{\psi(t)}]+\frac{1}{2}\big{(}C_{\psi(t)}P_{\psi(t)}+P_{\psi(t)}C_{\psi(t)}^{\dagger}(t)\big{)},}$ (15)
with $C_{\psi(t)}$ that can depend non-trivially on the current state of the realization $\psi(t)$. To emphasize the dependence on the state $\psi(t)$, we refer to the generalized RO as $\Psi$-ROQJ. Indeed, whenever $C_{\psi(t)}$ does not depend on $\psi(t)$, the $\Psi$-ROQJ reduces to the $R$-ROQJ of Eq. (13). The deterministic evolution is non-linear because of the state-dependence of the effective non-Hermitian Hamiltonian
${K_{\psi(t)}\coloneqq H(t)-\frac{i}{2}\Gamma(t)-\frac{i}{2}C_{\psi(t)}.}$ (16)
Since $C_{\psi(t)}$ appears only to applied to |$\ket{\psi(t)}$, it is possible to simplify the transformation by defining the unnormalized vector
|${\ket{\Phi_{\psi(t)}}\coloneqq C_{\psi(t)}\ket{\psi(t)}.}$ (17)
@@
${\begin{split}\Psi\text{-}R_{\psi(t)}=&\sum_{\alpha}\gamma_{\alpha}L_{\alpha}P_{\psi(t)}L^{\dagger}_{\alpha}\\ &+\frac{1}{2}\big{(}\ket{\Phi_{\psi(t)}}\bra{\psi(t)}+\ket{\psi(t)}\bra{\Phi_{\psi(t)}}\big{)}.\end{split}}$ (18)
The unraveling is obtained by considering jumps |$\ket{\psi(t)}\mapsto\ket{{\varphi_{\psi(t)}^{j}}}$ to the eigenstates of the RO, with probability $p^{j}_{\psi(t)}=\lambda_{\psi(t)}^{j}\,dt$, where $\lambda_{\psi(t)}^{j}$ is the corresponding eigenvalue. The deterministic evolution is, up to normalization, given by
|$\begin{split}\ket{\psi(t)}\mapsto&\ket{\tilde{\psi}^{\text{det}}(t+dt)}=(\mathbbm{1}-i{K_{\psi(t)}}\,dt)\ket{\psi(t)}\\ =&(\mathbbm{1}-iK(t)\,dt)\ket{\psi(t)}-\frac{dt}{2}\,{\ket{\Phi_{\psi(t)}}.}\end{split}$ (19)
In Appendix A, we show that averaging over all the realizations indeed reproduces the master equation (1). The key argument is that each state |$\ket{\psi_{i}(t)}$ evolves, on average, as it would according to (1), for any possible choice of $\Psi\text{-}R_{\psi_{i}(t)}$.
@@ IV Characterization and limitations
Figure 1: Examples of positivity domain $\mathscr{H}_{\mathcal{J}_{t}}$ (red), convex combination of projectors on elements of the positivity domain $\mathcal{P}(\mathscr{H}_{\mathcal{J}_{t}})$ (lighter red), and the time-evolved Bloch sphere $\Lambda_{t}(\mathcal{S}(\mathscr{H}))$ (green) for three non-P-divisible qubit dynamics at some fixed time $t$, showing the slice $y=0$ of the Bloch sphere. Left: no positive unraveling, since $\mathscr{H}_{\mathcal{J}_{t}}$ doesn’t contain any orthonormal basis. Middle: no positive unraveling, since $\Lambda_{t}\left(\mathcal{S}(\mathscr{H})\right)\not\subseteq\mathcal{P}(\mathscr{H}_{\mathcal{J}_{t}})$. Right: both necessary conditions hold, therefore a positive unraveling could exist.
@@ IV.1 Comparison with the $W$-ROQJ unravelings
By using the formalism of the $W$-ROQJ of Eq. (11), it is always possible to unravel any P-divisible dynamics by considering orthogonal jumps to the eigenstates of such operator. It is possible to obtain the same unravelings also with the $\Psi$-ROQJ formalism by imposing that $\psi(t)$ is an eigenstate of the RO: $\Psi\text{-}R_{\psi(t)}\ket{\psi(t)}=\lambda^{\text{det}}\ket{\psi(t)}$ for all $\psi(t)$, where $\lambda^{\text{det}}$ is the corresponding eigenvalue. This way, the other post-jump states |$\ket{\varphi^{j}_{\psi(t)}}$ are orthogonal to |$\ket{\psi(t)}$ and it can be shown that the corresponding eigenvalues are (see Appendix B)
${\lambda^{j}_{\psi(t)}=\sum_{\alpha}\gamma_{\alpha}\lvert\braket{\varphi^{j}_{\psi(t)}}{L_{\alpha}}{\psi(t)}\rvert^{2},}$ (20)
@@ IV.2 Necessary conditions for a positive unraveling
We now investigate when it possible to have a positive unraveling for dynamics breaking P-divisibility. This possibility drastically simplifies the simulations since, for positive unravelings, the different realizations do not depend on each other. For any non-P-divisible dynamics, there always exists some state $\psi$ and time $t$ such that ${\Psi\text{-}R_{\psi}}\not\geq 0$. In fact, from the condition for P-divisibility (2), there always exists a state $\psi_{\perp}$, orthogonal to $\psi$, such that
$\braket{\psi_{\perp}}{{\Psi\text{-}R_{\psi}}}{\psi_{\perp}}=\sum_{\alpha}\gamma_{\alpha}(t)\lvert\braket{\psi_{\perp}}{L_{\alpha}(t)}{\psi}\rvert^{2}<0.$ (21)
Notice that this condition does not depend on the particular transformation |$\ket{\Phi_{\psi}}$ present in (18), but only on $\mathcal{J}_{t}$. This fact, however, doesn’t necessarily limit the existence of a positive unraveling: it could still exist if one is able to describe the state $\rho(t)$ only using realizations |$\ket{\psi_{i}(t)}$ for which ${\Psi\text{-}R_{\psi_{i}(t)}}\geq 0$.
@@
where $\mathscr{H}$ is the system Hilbert space. From Eq. (2), it is evident that a dynamics is P-divisible if and only if $\mathscr{H}_{\mathcal{J}_{t}}=\mathscr{H}$ $\forall t$. At variance with the the positivity domain introduced e.g. in [53, 54], which refers to a subset of the set $\mathcal{L}(\mathscr{H})$ of linear operators on $\mathscr{H}$, the positivity domain defined in Eq. (22) is a subset of $\mathscr{H}$. However, $\mathscr{H}_{\mathcal{J}_{t}}$ is not a Hilbert space, and in particular it is not even a linear space.
A positive unraveling can exist only if it is possible to write any state as a convex combination $\rho(t)=\sum_{i}p_{i}(t){P_{\psi_{i}(t)}}$ using only states $\psi_{i}(t)\in\mathscr{H}_{\mathcal{J}_{t}}$ $\forall t$. Naturally, if $\mathscr{H}_{\mathcal{J}_{t}}=\emptyset$ for some time, it is not possible to have a positive unraveling. If we further assume invertibility of the dynamical map, then it is also not possible if $\mathscr{H}_{\mathcal{J}_{t}}$ is zero-measured: if, for any initial state, all stochastic realizations were inside a zero-measure set, then the whole set of quantum states $\mathcal{S}(\mathscr{H})$ would be mapped to a zero-measure set, thus breaking invertibility.
@@
These condition are depicted pictorially in Fig. 1. Interestingly, they do not depend on the unraveling but only on the dynamical map $\Lambda_{t}$. Therefore it is possible to rule out the possibility to have a positive unraveling for some dynamics by simply looking at $\Lambda_{t}$. On the other hand, as we will see in Sec. V.3 such conditions can be satisfied even by non P-divisible dynamics, which in fact do admit a positive unraveling